OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. answer choices: y – 2 = 3(x – 5) y + 8 = 3(x + 5) y + 5 = 3(x + 2) y – 5 = 3(x – 2)

OpenStudy (radar):

You have given the slope (3) and a point P1(2,5). This is all you need to calculate the slope intercept form for this line. Since slope is equal to the change in y for a corresponding in x you can do this:

OpenStudy (anonymous):

I thought it was y + 5 = 3(x + 2)

OpenStudy (anonymous):

I'm not sure though, How would i do it

OpenStudy (radar):

(5-Y)/2-x) = 3 3(2-x)=5-y 6-3x=5-y transposing terms y=3x -6+5 y=3x-1

OpenStudy (radar):

The last one can be reduced to this. y-5=3(x-2)

OpenStudy (anonymous):

Alright thank youuu @radar

OpenStudy (radar):

You're welcome.